Let X be a continuous r.v with density function
f(x) = ( |x|
10
, for − 2 < x < 4
0, elsewhere
. Calculate the expected value of X.
1
Expert's answer
2020-05-12T19:04:02-0400

E(x)=∫ x*f(x)dx=∫ 4-2x*|x|/10=(∫ 0-2x*(-x)dx)/10+(∫ 40x2dx)/10=-4/15+32/15=28/15=1.86